# Convexified Message-Passing Graph Neural Networks

### v1

@@ 4.5 Convergence and Generalization Error
We now analyze Algorithm 2’s generalization error, focusing on models with a single hidden layer ($L=1$) in the binary classification case where the output dimension is $G=1$. Our main result can be easily extended to the multi-class case by a standard one-versus-all reduction to the binary case. In case of multiple hidden layers, our analysis then applies to each hidden layer separately. Specifically, by an approach similar to Zhang et al. [59], we prove that the generalization error of Algorithm 2 converges to the optimal generalization error of a GCN, as it is bounded above by this optimal error plus an additive term that decays polynomially to zero with the sample size. For any $\ell\in[L]$ and $k\in[K]$, we use the random kernel matrix $\mathcal{K}_{\ell}^{k}(\mathcal{X})\in\mathbb{R}^{n\times n}$ induced by a graph signal $\mathcal{X}\in\mathbb{R}^{n\times F}$ drawn randomly from the population, i.e., the $(i,i^{\prime})$-th entry of $\mathcal{K}_{\ell}^{k}(\mathcal{X})$ is $\kappa_{\ell}(\mathbf{z}_{i}^{k}(\mathcal{X}),\mathbf{z}_{i^{\prime}}^{k}(\mathcal{X}))$, where $\mathbf{z}_{i}^{k}(\mathcal{X}_{\ell-1}):=[\mathcal{S}^{k}\mathcal{X}_{\ell-1}]_{i}$ aggregates information at node $i$ from $k$-hops away.
@@
Consider models with a single hidden layer for binary classification (i.e., $L=G=1$). Let $J(\cdot,\mathbf{y})$ be $M$-Lipchitz continuous for any fixed $\mathbf{y}\in[G]^{n}$ and let $\kappa_{L}$ be either the inverse polynomial kernel or the Gaussian RBF kernel (Recall Appendices B.1-B.2). For any valid activation function $\sigma$ (Recall Appendix B.3), there is a constant radius $R_{L}>0$ such that the expected generalization error is at most:
$\mathbb{E}_{\mathcal{X},\mathbf{y}}[J(\hat{\Psi}^{\hat{\mathcal{A}}_{L}}(\mathcal{X}),\mathbf{y})]\leq\inf_{\Phi\in\mathcal{F}_{\text{gcn}}}\mathbb{E}_{\mathcal{X},\mathbf{y}}[J(\Phi(\mathcal{X}),\mathbf{y})]+\frac{Const}{\sqrt{m}},$
@@
(Sketch) The full proof in Appendix F comprises of two steps: When $L=G=1$, we first consider a wealthier function class, denoted as $\mathcal{F}_{\text{cgcn}}$, which is proven to contain the class of GCNs $\mathcal{F}_{\text{gcn}}$ from (2). The richer class relaxes the class of GCNs in two aspects: (1) The graph filters are relaxed to belong to the RKHS induced by the kernel, and (2) the $\ell_{2}$-norm constraints in (2) are substituted by a single constraint on the Hilbert norm induced by the kernel. Though Algorithm 2 never explicitly optimizes over the function class $\mathcal{F}_{\text{cgcn}}$, we prove that the predictor $\hat{\Psi}^{\hat{\mathcal{A}}_{L}}$ generated at the last layer is an empirical risk minimizer in $\mathcal{F}_{\text{cgcn}}$. Then, we prove that $\mathcal{F}_{\text{cgcn}}$ is not "too big" by upper bounding its Rademacher complexity, a key concept in empirical process theory that can be used to bound the generalization error in our ERM problem (see, e.g., [2] for a brief overview on Rademacher complexity). Combining this upper bound with standard Rademacher complexity theory [2], we infer that the generalization error of $\hat{\Psi}^{\hat{\mathcal{A}}_{L}}$ converges to the best generalization error of $\mathcal{F}_{\text{cgcn}}$. Since $\mathcal{F}_{\text{gcn}}\subset\mathcal{F}_{\text{cgcn}}$, the latter error is bounded by that of GCNs, completing the proof. ∎
@@ 5 Experimental Evaluation
Datasets. We evaluate our framework through extensive experiments on several benchmark graph classification datasets [31], including the four bioinformatics datasets MUTAG, PROTEINS, PTC-MR, NCI1 and the chemical compounds dataset Mutagenicity. Following many prior works [29, 53], we randomly split each dataset into three parts: 80% as training set, 10% as validation set and the remaining 10% as test set. The statistics of these datasets appear in Appendix G.1.
@@
Baselines. We apply our convexification procedure to the message-passing mechanisms of popular GNNs and hybrid graph transformers, creating convex counterpartsc. Our framework is integrated into a wide range of leading models, including GCN [7, 24], GAT [45], GATv2 [4], GIN [51], GraphSAGE [20], ResGatedGCN [3], as well as the hybrid transformers GraphGPS [36] and GraphTrans [40, 49]. All models share the same architecture design for fair comparison; see Appendix G.2 for details.
Optimization. All models are trained over $200$ epochs with the Adam optimizer [23] using the cross entropy loss, a starting learning rate of $10^{-3}$, a minimum learning rate of $10^{-6}$ and a weight decay of $10^{-5}$. We use a ReduceLROnPlateau scheduler, which cuts the learning rate in half with a patience of $20$ epochs, and training ends when we reach the minimum learning rate.
Results and Analysis. Table 1 presents our main empirical results, comparing two-layer convex models against two-layer and six-layer non-convex models. Impressively, in most cases, the two-layer convex variants substantially surpass their non-convex counterparts by 10–40% accuracy, showcasing the practical strength of our framework alongside its theoretical benefits. Even when improvements are modest, our convex models match or slightly surpass their non-convex versions, underscoring the robustness and broad applicability of our approach. Notable examples are GATv2 and GraphGPS, whose convex versions significantly outperform their non-convex ones in 4 out of 5 datasets, with slight improvements on Mutagenicity. Our convex models’ superiority over the hybrid transformer GraphGPS, which combines message-passing with self-attention, highlights the power of convexified GNNs over non-convex ones, even when enhanced by self-attention. To visualize these trends more holistically, we also include additional plots in Appendix H.
Over-parameterization is known to aid gradient-based methods to reach global minima that generalize well in non-convex models [1, 11, 60], making it a common practice for obtaining strong performance. Our results challenge this approach. We show that our shallow convex models often outperform their more over-parameterized non-convex counterparts, obtaining better results with fewer layers and parameters. In effect, we beat them twice: in performance and in resource efficiency. These gains are especially notable in small-data regimes, which are prone to overfitting, where we suspect that local optima obtained by non-convex models may hinder generalization. In contrast, convex models avoid such pitfalls, offering more stable and reliable learning. Our findings show that rich representations can be attained without over-parameterization by exploiting convexity in shallow models.
@@
Table 1: Accuracy (%±$\%\pm\sigma$) of two-layer convex versions of base models against their two-layer and six-layer non-convex counterparts on various graph classification datasets over $4$ runs, each with different seeds. The datasets are ordered from left to right in ascending order of size. Convex variants start with ’C’. A two-layer model is marked by (2L), while a six-layer one is marked by (6L). Results of convex models surpassing their non-convex variants by a large gap ($\sim$10-40%) are in dark green and modest improvements in light green.
@@
CGCN (2L) $\mathbf{71.2\pm 2.1}$ $\mathbf{59.7\pm 4.2}$ $\mathbf{70.9\pm 2.4}$ $\mathbf{62.3\pm 1.1}$ $\mathbf{61.9\pm 2.7}$
GCN (2L) $55.0\pm 22.0$ $48.5\pm 10.4$ $59.5\pm 4.1$ $50.8\pm 1.7$ $54.7\pm 1.1$
GCN (6L) $61.2\pm 21.6$ $55.0\pm 15.8$ $58.7\pm 6.1$ $50.8\pm 1.7$ $54.7\pm 1.9$
CGIN (2L) $\mathbf{83.2\pm 7.3}$ $\mathbf{63.2\pm 4.7}$ $\mathbf{70.1\pm 2.6}$ $\mathbf{63.6\pm 2.3}$ $\mathbf{59.1\pm 1.8}$
GIN (2L) $60.0\pm 17.6$ $45.0\pm 9.5$ $50.4\pm 12.0$ $55.5\pm 1.4$ $54.8\pm 2.0$
GIN (6L) $61.2\pm 1.6$ $48.5\pm 13.4$ $61.6\pm 1.2$ $60.8\pm 2.8$ $55.8\pm 2.0$
CGAT (2L) $\mathbf{72.5\pm 4.7}$ $\mathbf{59.7\pm 6.0}$ $\mathbf{71.8\pm 2.5}$ $\mathbf{64.6\pm 0.0}$ $\mathbf{58.8\pm 3.6}$
GAT (2L) $65.0\pm 9.3$ $51.4\pm 15.7$ $60.0\pm 3.3$ $51.9\pm 2.2$ $54.8\pm 2.0$
GAT (6L) $68.7\pm 4.1$ $54.2\pm 17.2$ $63.1\pm 1.8$ $52.3\pm 2.9$ $54.8\pm 2.0$
CGATv2 (2L) $\mathbf{68.7\pm 4.1}$ $\mathbf{59.0\pm 5.2}$ $\mathbf{71.3\pm 0.6}$ $\mathbf{63.6\pm 1.1}$ $\mathbf{62.3\pm 4.2}$
GATv2 (2L) $55.0\pm 19.0$ $49.2\pm 15.3$ $61.6\pm 1.6$ $49.5\pm 2.0$ $55.9\pm 1.4$
GATv2 (6L) $57.5\pm 16.0$ $49.2\pm 15.3$ $63.3\pm 2.1$ $49.5\pm 2.0$ $59.6\pm 6.2$
CResGatedGCN (2L) $\mathbf{68.7\pm 4.1}$ $\mathbf{60.4\pm 5.8}$ $\mathbf{72.5\pm 1.0}$ $\mathbf{62.4\pm 3.2}$ $\mathbf{59.6\pm 3.7}$
ResGatedGCN (2L) $57.0\pm 16.7$ $50.7\pm 17.7$ $47.9\pm 9.9$ $56.8\pm 6.6$ $56.1\pm 2.1$
ResGatedGCN (6L) $65.0\pm 15.8$ $50.7\pm 7.1$ $55.8\pm 8.7$ $61.1\pm 3.1$ $56.1\pm 2.1$
CGraphSAGE (2L) $\mathbf{68.7\pm 4.1}$ $\mathbf{65.3\pm 11.5}$ $\mathbf{73.8\pm 3.7}$ $\mathbf{63.7\pm 1.0}$ $\mathbf{66.3\pm 5.3}$
GraphSAGE (2L) $57.4\pm 16.0$ $52.8\pm 7.4$ $62.5\pm 0.8$ $53.1\pm 6.0$ $58.1\pm 3.6$
GraphSAGE (6L) $51.2\pm 1.9$ $55.7\pm 12.5$ $63.1\pm 1.8$ $53.2\pm 6.1$ $60.0\pm 3.3$
CGraphTrans (2L) $\mathbf{68.7\pm 4.1}$ $\mathbf{63.3\pm 10.8}$ $\mathbf{70.7\pm 3.1}$ $\mathbf{63.1\pm 2.8}$ $\mathbf{63.2\pm 4.7}$
GraphTrans (2L) $57.4\pm 16.0$ $40.4\pm 8.0$ $60.7\pm 2.6$ $55.0\pm 3.7$ $53.6\pm 1.3$
GraphTrans (6L) $67.5\pm 7.4$ $49.2\pm 12.3$ $61.6\pm 1.2$ $61.5\pm 2.5$ $59.9\pm 4.6$
CGraphGPS (2L) $\mathbf{83.7\pm 2.1}$ $\mathbf{65.5\pm 9.3}$ $\mathbf{81.6\pm 7.2}$ $\mathbf{60.1\pm 6.1}$ $\mathbf{59.4\pm 4.1}$
GraphGPS (2L) $50.9\pm 19.6$ $49.2\pm 20.1$ $38.0\pm 1.9$ $50.8\pm 2.6$ $41.4\pm 8.7$
GraphGPS (6L) $52.5\pm 2.1$ $46.4\pm 2.3$ $38.8\pm 1.5$ $49.4\pm 1.9$ $55.1\pm 5.5$
6 Conclusions and Future Work

### v2

@@ 4.5 Convergence and Generalization Error
We now analyze Algorithm 2’s generalization error, focusing on two-layer models ($L=1$) in the binary classification case where the output dimension is $G=2$. Our main result can be easily extended to the multi-class case by a standard one-versus-all reduction to the binary case. In case of multiple hidden layers, our analysis then applies to each hidden layer separately. However, a global depth-wide bound is an important future direction.
Specifically, by an approach similar to Zhang et al. (2017), we prove that the generalization error of Algorithm 2 converges to the optimal generalization error of a GCN, as it is bounded above by this optimal error plus an additive term that decays polynomially to zero with the sample size. For any $\ell\in[L]$ and $k\in[K]$, we use the random kernel matrix $\mathcal{K}_{\ell}^{k}(\mathcal{X})\in\mathbb{R}^{n\times n}$ induced by a graph signal $\mathcal{X}\in\mathbb{R}^{n\times F}$ drawn randomly from the population, i.e., the $(i,i^{\prime})$-th entry of $\mathcal{K}_{\ell}^{k}(\mathcal{X})$ is $\kappa_{\ell}(\mathbf{z}_{i}^{k}(\mathcal{X}),\mathbf{z}_{i^{\prime}}^{k}(\mathcal{X}))$, where $\mathbf{z}_{i}^{k}(\mathcal{X}_{\ell-1}):=[\mathcal{S}^{k}\mathcal{X}_{\ell-1}]_{i}$ aggregates information at node $i$ from $k$-hops away.
@@
Consider two-layer models in the binary classification case (i.e., $L=1,G=2$). Let $J(\cdot,\mathbf{y})$ be $M$-Lipchitz continuous for any fixed $\mathbf{y}\in[G]^{n}$ and let $\kappa_{L}$ be either the inverse polynomial kernel or the Gaussian RBF kernel (Recall Appendices C.1-C.2). For any valid activation function $\sigma$ (Recall Appendix C.3), there is a constant radius $R_{L}>0$ such that the expected generalization error is at most:
$\mathbb{E}_{\mathcal{X},\mathbf{y}}[J(\hat{\Psi}^{\hat{\mathcal{A}}_{L}}(\mathcal{X}),\mathbf{y})]$ $\leq\inf_{\Phi\in\mathcal{F}_{\text{gcn}}}\mathbb{E}_{\mathcal{X},\mathbf{y}}[J(\Phi(\mathcal{X}),\mathbf{y})]$
$\quad+\frac{Const}{\sqrt{m}},$
@@
(Sketch) The full proof in Appendix G comprises of two steps: When $L=1,G=2$, we first consider a wealthier function class, denoted as $\mathcal{F}_{\text{cgcn}}$, which is proven to contain the class of GCNs $\mathcal{F}_{\text{gcn}}$ from (2). The richer class relaxes the class of GCNs in two aspects: (1) The graph filters are relaxed to belong to the RKHS induced by the kernel, and (2) the $\ell_{2}$-norm constraints in (2) are substituted by a single constraint on the Hilbert norm induced by the kernel. Though Algorithm 2 never explicitly optimizes over the function class $\mathcal{F}_{\text{cgcn}}$, we prove that the predictor $\hat{\Psi}^{\hat{\mathcal{A}}_{L}}$ generated at the last layer is an empirical risk minimizer in $\mathcal{F}_{\text{cgcn}}$. Then, we prove that $\mathcal{F}_{\text{cgcn}}$ is not ”too big” by upper bounding its Rademacher complexity, a key concept in empirical process theory that can be used to bound the generalization error in our ERM problem (see, e.g., (Bartlett and Mendelson, 2002) for a brief overview on Rademacher complexity). Combining this upper bound with standard Rademacher complexity theory (Bartlett and Mendelson, 2002), we infer that the generalization error of $\hat{\Psi}^{\hat{\mathcal{A}}_{L}}$ converges to the best generalization error of $\mathcal{F}_{\text{cgcn}}$. Since $\mathcal{F}_{\text{gcn}}\subset\mathcal{F}_{\text{cgcn}}$, the latter error is bounded by that of GCNs, completing the proof. ∎
@@ 5 EXPERIMENTAL EVALUATION
Datasets. We evaluate our framework through extensive experiments on several benchmark graph classification datasets (Morris et al., 2020), including the four bioinformatics datasets MUTAG, PROTEINS, PTC-MR, NCI1 and the chemical compounds dataset Mutagenicity. We also evaluated our models on the ogbg-molhiv molecular property prediction dataset from the Open Graph Benchmark (Hu et al., 2020), originally sourced from MoleculeNet (Wu et al., 2018) and among its largest benchmarks. Following many prior works (Ma et al., 2019; Ying et al., 2018a), we randomly split each dataset into three parts: 80% as training set, 10% as validation set and the remaining 10% as test set. The statistics of the above datasets appear in Appendix H.1.
@@
Baselines. We apply our convexification procedure to the message-passing mechanisms of popular GNNs and hybrid graph transformers, creating convex counterparts. Our framework is integrated into a wide range of leading models, including GCN (Defferrard et al., 2016; Kipf and Welling, 2017), GAT (Veličković et al., 2018), GATv2 (Brody et al., 2022), GIN (Xu et al., 2019), GraphSAGE (Hamilton et al., 2017), ResGatedGCN (Bresson and Laurent, 2017), as well as the hybrid transformers GraphGPS (Rampášek et al., 2022) and GraphTrans (Shi et al., 2021; Wu et al., 2021). In Appendix I.5, we include additional experiments for DirGNN (Rossi et al., 2024). All models share the same architecture design for fair comparison; see Appendix H.2 for more details. Our code is available at (Cohen et al., 2026).
We would like to clarify that many recent graph learning papers in influential venues continue using the same baselines or older ones (see, e.g., (Armgaan et al., 2025; Yehudai et al., 2025; Hordan et al., 2025; Ma et al., 2023; Chen et al., 2023a)). Further, our baselines are still listed in leaderboards of modern datasets like Open Graph Benchmark (OGB) (Hu et al., 2020). Thus, our baselines match what current peer-reviewed GNN work uses for molecular datasets and graph classification, especially for message-passing comparisons. We also clarify that our goal is parallel: to compare convexified GNNs to their own non-convex counterparts, not to propose a new SOTA model.
@@
Table 1: Accuracy ($\%\pm\sigma$) of 22-layer convex versions of base models against their 22- and 66-layer non-convex counterparts on various graph classification datasets over 44 runs, each with different seeds. Datasets are ordered from left to right in ascending order of size. Convex variants start with ’C’. A 22-layer model is marked by (2L), while a 66-layer one by (6L). Results of convex models surpassing their non-convex variants by a large gap ($\sim$10–40%) are in bold and modest improvements are underlined.
@@
CGCN (2L) $\mathbf{85.0\pm 3.5}$ $67.6\pm 5.8$} $\mathbf{75.0\pm 3.2}$ $\mathbf{77.1\pm 2.8}$ $80.6\pm 3.3$}
GCN (2L) $50.0\pm 20.3$ $48.5\pm 10.4$ $59.5\pm 4.1$ $61.0\pm 4.0$ $69.7\pm 2.4$
GCN (6L) $65.0\pm 19.6$ $63.0\pm 11.8$ $58.7\pm 6.1$ $65.0\pm 5.9$ $72.7\pm 1.6$
CGIN (2L) $\mathbf{82.5\pm 5.5}$ $\mathbf{66.4\pm 3.4}$ $\mathbf{77.7\pm 1.8}$ $74.8\pm 9.4$} $\mathbf{77.9\pm 4.5}$
GIN (2L) $50.0\pm 11.7$ $49.3\pm 10.0$ $50.4\pm 12.0$ $61.0\pm 1.9$ $67.8\pm 4.6$
GIN (6L) $61.2\pm 14.3$ $51.9\pm 14.5$ $1.6\pm 1.2 66$$6.3\pm 2.4$ $69.1\pm 3.2$
CGAT (2L) $\mathbf{83.7\pm 5.4}$ $\mathbf{70.1\pm 9.9}$ $\mathbf{72.1\pm 4.7}$ $68.5\pm 4.7$} $75.3\pm 2.6$}
GAT (2L) $62.0\pm 22.7$ $58.7\pm 18.3$ $60.0\pm 3.3$ $67.1\pm 3.9$ $69.3\pm 3.0$
GAT (6L) $75.0\pm 16.2$ $61.0\pm 17.6$ $63.1\pm 1.8$ $67.3\pm 2.6$ $72.8\pm 2.0$
CGATv2 (2L) $\mathbf{87.5\pm 2.5}$ $\mathbf{69.4\pm 7.5}$ $\mathbf{73.4\pm 1.7}$ $66.2\pm 1.5$} $80.6\pm 1.0$}
GATv2 (2L) $68.7\pm 4.1$ $56.3\pm 13.6$ $61.6\pm 1.6$ $62.2\pm 6.2$ $70.1\pm 2.7$
GATv2 (6L) $70.0\pm 7.9$ $59.2\pm 17.4$ $63.3\pm 2.1$ $63.5\pm 4.0$ $74.9\pm 1.9$
CResGatedGCN (2L) $\mathbf{83.7\pm 5.4}$ $\mathbf{73.0\pm 5.0}$ $\mathbf{78.5\pm 2.2}$ $69.8\pm 3.2$} $79.1\pm 3.5$}
ResGatedGCN (2L) $43.7\pm 17.4$ $50.7\pm 7.1$ $47.9\pm 9.9$ $67.2\pm 2.7$ $70.6\pm 0.9$
ResGatedGCN (6L) $5.0\pm 12.2 55$$4\pm 20.5 55.$$5.8\pm 8.7$ $67.3\pm 2.5$ $71.5\pm 1.6$
CGraphSAGE (2L) $\mathbf{82.5\pm 2.4}$ $\mathbf{72.3\pm 6.1}$ $\mathbf{73.6\pm 3.1}$ $69.3\pm 2.8$} $\mathbf{80.1\pm 3.6}$
GraphSAGE (2L) $65.0\pm 16.5$ $56.2\pm 17.0$ $62.5\pm 0.8$ $61.4\pm 4.1$ $70.6\pm 1.8$
GraphSAGE (6L) $55.0\pm 12.7$ $61.5\pm 8.9$ $63.1\pm 1.8$ $64.7\pm 2.0$ $70.8\pm 2.3$
CGraphTrans (2L) $\mathbf{81.2\pm 4.1}$ $\mathbf{68.1\pm 8.3}$ $\mathbf{75.4\pm 3.2}$ $68.9\pm 2.5$} $73.8\pm 3.4$}
GraphTrans (2L) $62.5\pm 16.7$ $55.3\pm 13.8$ $60.7\pm 2.6$ $68.9\pm 0.8$ $68.0\pm 4.1$
GraphTrans (6L) $57.5\pm 19.5$ $4\pm 16.9 61.6±1.$$6\pm 1.2$ $70.0\pm 4.3$ $72.1\pm 3.8$
CGraphGPS (2L) $\mathbf{83.7\pm 2.1}$ $\mathbf{65.5\pm 9.3}$ $\mathbf{81.6\pm 7.2}$ $\mathbf{60.1\pm 6.1}$ $59.4\pm 4.1$}
GraphGPS (2L) $50.9\pm 19.6$ $49.2\pm 20.1$ $42.0\pm 1.9$ $50.8\pm 2.6$ $41.4\pm 8.7$
GraphGPS (6L) $52.5\pm 2.1$ $46.4\pm 2.3$ $48.8\pm 1.5$ $49.4\pm 1.9$ $55.1\pm 5.5$
Results and Analysis. Table 1 presents our main empirical results for the graph classification benchmarks, comparing two-layer convex models against two-layer and six-layer non-convex models. On the ogbg-molhiv dataset, our convex models consistently obtained around 98% accuracy, surpassing their non-convex counterparts, which reached about 96%. Impressively, in most cases, the two-layer convex variants substantially surpass their non-convex counterparts by 10–40% accuracy, showcasing the practical strength of our framework alongside its theoretical benefits. Even when improvements are modest, our convex models match or slightly surpass their non-convex versions, underscoring the robustness and broad applicability of our approach. Notable examples are GATv2 and GraphGPS, whose convex versions significantly outperform their non-convex ones in 4 out of 5 datasets, with slight improvements on Mutagenicity. Our convex models’ superiority over the hybrid transformer GraphGPS, which combines message-passing with self-attention, highlights the power of convexified GNNs over non-convex ones, even when enhanced by self-attention. In Appendix I, we visualize these trends more holistically.
Over-parameterization is known to aid gradient-based methods to reach global minima that generalize well in non-convex models (Allen-Zhu et al., 2019; Du et al., 2018; Zou et al., 2020), making it a common practice for obtaining strong performance. Our results challenge this approach, showing that our shallow convex models often surpass their more over-parameterized non-convex counterparts and obtain better results with fewer layers and parameters. Namely, we surpass them in both performance and model compactness. These gains are most notable in small-data regimes, which are prone to overfitting, where we suspect that local optima obtained by non-convex models may hinder generalization. However, convex models avoid such pitfalls, offering more stable and reliable learning. Our results show that rich representations can be attained without over-parameterization by exploiting convexity in shallow models.
In Appendix I.4, we give more results for deeper convex models, depicting how accuracy can be improved as depth increases. This directly addresses how our convexified models handle over-smoothing: our results in I.4 show that the accuracy of convex models can further improve with additional layers, which is consistent with our theoretical guarantees in Theorem 4.7. Hence, this illustrates that errors remain controlled as depth increases, offering strong empirical support for the practical effectiveness of multi-layer CGNNs, which is crucial because state-of-the-art GNN applications often rely on deeper architectures. Intuitively, convexified models handle over-smoothing well thanks to the nuclear norm relaxation used by our model. Specifically, the nuclear norm projection reduces the effective capacity of filters and, in practice, suppresses small singular-value directions that often correspond to high-frequency/noisy modes over the graph. This prevents representation collapse in shallow models and empirically reduces over-smoothing in deeper convex models. See Appendix I.6 for more details.
6 CONCLUSION